OpenStudy (anonymous):

How do I rewrite cos3x with only sinx and cosx?

OpenStudy (anonymous):

cos(3x) = cos(2x + x) = cos(2x)cos(x) - sin(2x)sin(x)

OpenStudy (anonymous):

Use the sum / difference identities Cos(x+y)= cos(x)cos(y) - sin(x)sin(y)